sample of impure Magnesium is reacted with diluted Sulphuric Acid to give the respective salt and Hydrogen.If 1 gram of impure sample gave 298.6 cc of Hydrogen at STP.Calculate the % purity of the sample (MAgnesium = 24,Hydrogen = 1)


Grade:

1 Answers

Vikas TU
10 years ago

Magnesium is reacted with diluted Sulphuric Acid gives the following reaction:

            

         Moles of H2 at STP = (298.6)/(22.4 x 1000) = 0.0133  

         Mg        +         H2SO4  ------->   MgSO4            +              H2

                                                                                                      0.0133

Thus, stoich. coeff. are equal hence moles of Mg are same as H2 .i.e = 0.0133

Now, Mg in gm = moles x molar mass of Mg

                           = 0.0133 x 24

                           =0.3192 gm

Now percentage of purity of the sample is = 0.3192/1 x 100

                                                                            = 31.9 or 32 % (approx)

 

 

Plz Approve!

 

Plz Approve!

 

Plz Approve!